The table below shows the values for several different components of GDP Value (billions of dollars) 266 Personal consumption expenditures Net foreign factor income Transter payments Rents Consumption of fixed capital (depreciation Statistical discrepancy Social Security contributions Interest Proprietors income Net exports Dividends Compensation of employees Taxes on production and imports Undistributed corporate profits Personal taxes Corporate income taxes Corporate profits 15 12 10 23 12 16 263 16 46 61 Government purchases Net private domestic investment Personal saving 10 5 ImportsInstructions: Enter your answers as whole numbers a. What is the value of exports? billion b. What is the value of national income? billion c. What is the value of GDP? billion d. What is the value of personal income? billion
Answer #1

(a) Calculate the value of exports: Net Exports = Exports - Imports Exports = Net Exports+Imports = 12 + 5 = 17 Therefore, th(b) Calculate NDP: NDP = GDP - Consumption of Fixed Capital =388-23 = $365 Calculate National Income: NI = NDP - Statistical(d) Calculate Personal Income: (NI - Corporate Income Tax - Undistributed Corporate Profits Personal Income= 1-Social Securit